Use inductive reasoning to predict the most probable next number in each of the following lists. b. 2, 5, 10, 17, 26, (?)


1
Expert's answer
2021-09-20T23:20:16-0400

Study the sequence,

We realize that we are supposed to add odd numbers starting from 3 to the next.

2

2+3=5

5+5=10

10+7=17

17+9=26

26+11=37


Most probable next number =37